### v1

# WaveGNN: Modeling Irregular Multivariate Time Series for Accurate Predictions

Abstract
Accurately modeling and analyzing time series data is crucial for downstream applications across various fields, including healthcare, finance, astronomy, and epidemiology. However, real-world time series often exhibit irregularities such as misaligned timestamps, missing entries, and variable sampling rates, complicating their analysis. Existing approaches often rely on imputation, which can introduce biases. A few approaches that directly model irregularity tend to focus exclusively on either capturing intra-series patterns or inter-series relationships, missing the benefits of integrating both. To this end, we present WaveGNN, a novel framework designed to directly (i.e., no imputation) embed irregularly sampled multivariate time series data for accurate predictions. WaveGNN utilizes a Transformer-based encoder to capture intra-series patterns by directly encoding the temporal dynamics of each time series. To capture inter-series relationships, WaveGNN uses a dynamic graph neural network model, where each node represents a sensor, and the edges capture the long- and short-term relationships between them. Our experimental results on real-world healthcare datasets demonstrate that WaveGNN consistently outperforms existing state-of-the-art methods, with an average relative improvement of $14.7\%$ in F1-score when compared to the second-best baseline in cases with extreme sparsity. Our ablation studies reveal that both intra-series and inter-series modeling significantly contribute to this notable improvement.
1 Introduction
The widespread use of sensors across various sectors, including healthcare, finance, astronomy, and urban applications [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], has led to a significant increase in the collection of time series data, ranging from univariate to multivariate types, where multiple measures are recorded simultaneously. Effectively modeling and analyzing these time series datasets is crucial for downstream applications such as monitoring patient health, optimizing treatments, and managing disease outbreaks, where accurate and timely data interpretation can significantly impact outcomes. In real-world applications, though, time series often exhibit irregularities, frequently caused by sensor malfunctions, different sampling rates across sensors, and cost-saving measures [19]. This variability in data collection renders traditional machine learning (ML) approaches, which typically rely on a fixed number of regularly sampled observations across all sensors, ineffective for accurate predictions [20, 19].
Modeling irregularly sampled time series can be complex due to several factors: measurements from different sensors may not align, data may be missing, and each measurement sequence may vary in size [19, 21, 22]. To address these challenges, recent developments have concentrated on creating specialized algorithms and model architectures designed to effectively manage these irregularities [23].
One category of approaches tackles irregularity by filling in the missing values, thereby converting the datasets into regularly sampled time series. These imputation techniques vary from traditional hand-crafted imputation methods [24] to more advanced learned interpolation methods [25, 21]. Once the missing values are filled, the resulting regular time series are utilized to perform the downstream predictions [26, 27, 28]. While these techniques address data gaps, they introduce biases and distortions in the data distribution. Furthermore, the absence of measurements can itself provide valuable insights; for instance, periodically missing heart rate data might indicate that a patient underwent a medical procedure during which the monitoring device was removed.
Another category of approaches focuses on directly learning from the irregularities by designing specialized model architectures [21, 29, 19, 20]. For instance, the GRU-D model incorporates a decaying mechanism within gated recurrent units (GRUs) to better handle irregular sampling intervals [29], while mTAND utilizes multi-time attention to analyze non-uniform measurements effectively [21]. These techniques predominantly capture the correlations within a single sensor’s measurements over time, focusing on intra-series dynamics. Meanwhile, other methods have been proposed to capture inter-series dependencies, modeling explicitly the relationships across sensors within the dataset [19].
Despite these advancements, most approaches tend to specialize in either capturing intra-series dynamics or inter-series relationships, but rarely both simultaneously. Yet, effectively modeling both types of relationships must lead to more accurate predictions and deeper insights. For example, understanding both the individual progression of patients’ vital signs (e.g., heart rate) and how they interact with one another (e.g., the co-dependencies between heart rate and blood pressure over time) can provide a holistic view of a patient’s condition, potentially improving diagnostic accuracy and treatment outcomes. Additionally, in scenarios where one or more sensor observations are missing, irregularity can be compensated by exploiting insights among the correlated sensors.
Inspired by these insights, we introduce WaveGNN, a novel framework designed to directly (without imputation) embed irregularly sampled multivariate time series datasets for accurate predictions. WaveGNN captures both the intra-series patterns and the inter-series relationships simultaneously, with two main components. To effectively capture the temporal dependencies of each individual time series, the first component is a Transformer-based encoder. This encoder incorporates a masked attention mechanism that excludes missing data observations during processing, along with a relative timestamp encoding mechanism to capture the time intervals between consecutive observations. Given that more recent observations can be more indicative of the final prediction, a decay mechanism is also incorporated within the encoder to prioritize newer observations. To adequately capture inter-sensor dependencies, the second component of WaveGNN is a dynamic graph neural network, where each node corresponds to a sensor in the dataset, and the edges between nodes denote the correlations between these sensors. These edges are dynamically inferred based on sequence embeddings generated by the transformer-based encoder during the time window and learned global node embeddings shared across all samples and all time windows. Consequently, WaveGNN effectively captures both the dynamic short-term relationships observed within each window and the long-term relationships across the entire dataset, effectively mitigating the impact of missing data by leveraging inter-sensor information.
We evaluate WaveGNN across four real-world healthcare datasets, each associated with distinct classification tasks. Our experiments show that WaveGNN consistently outperforms both imputation-based and specialized state-of-the-art baselines by up to 10.2% in AUPRC and 3.4% in accuracy. Even in real-world scenarios where entire sets of sensor observations are missing, WaveGNN maintains high performance, outperforming baselines with an average relative improvement of 14.7% in F1 score on the challenging PAM dataset, which involves an 8-class classification task. Our ablation study reveals that each proposed component contributes to the overall effectiveness of WaveGNN, with inter-series and intra-series dependencies showing the most significant performance drops when excluded, leading to F1 score reductions of 33.8% and 13.2%, respectively.
2 Related Work
Approaches for handling irregularities in multivariate time series vary widely. Temporal Discretization-based Embedding (TDE) methods transform irregular time series into regular forms through imputation [26]. Such approaches enable the use of standard deep learning models but inherently introduce biases and distort data distributions [27, 28, 24]. To address their shortcomings, approaches such as IP-Net and DGM2 proposed using kernel functions to interpolate data against reference points to achieve temporal alignment [25, 30]. Shukla et al. further proposed a time attention mechanism with time embeddings to learn interpolation representations [21], while Zhang et. al [26] introduced a hybrid model that combines different TDE methods and integrated hand-crafted imputation embeddings into learned interpolation embeddings to improve medical predictions [26]. These approaches still fundamentally rely on imputation, retaining the inherent drawbacks of TDE methods.
Another group of approaches model irregularity directly through specialized algorithms or model architecture designs. To that end, GRU-D employs decaying mechanisms within GRUs to handle the irregular sampling of the data [29]. SeFT applies set function learning to time series classification, treating the series as a set of observations rather than a sequence [20]. Additionally, Raindrop employs graph neural networks to conceptualize multivariate time series as separate sensor graphs, capturing complex inter-variable relationships [19]. While these approaches avoid the TDE drawbacks, they often require extensive customization and typically focus on capturing either intra-variable dynamics or inter-variable dependencies, but not both, potentially overlooking essential data patterns. In contrast, WaveGNN introduces a more integrated approach, utilizing a standard transformer architecture with masking and time encoding to effectively model intra-series relationships and a graph neural network architecture to capture short and long-term inter-variable dependencies simultaneously. We evaluate WaveGNN against the aforementioned approaches in Section 5.2.
Graph-based approaches have been extensively applied to model multivariate time series datasets. As an example of dynamic graph-based approaches, BysGNN creates a dynamic multi-context graph for forecasting visits to points of interest (POIs) based on spatial, temporal, semantic, and taxonomic contexts [6]. Similarly, NeuroGNN, captures multi-context correlations in EEG data for seizure classification [4]. These approaches are designed for regular time series data and rely on having external context information to accurately infer inter-series relationships. In contrast, WaveGNN models the inter-series relationships given irregular multivariate time series without requiring additional context.
Figure 1: Overall WaveGNN Framework
3 Preliminaries
Definition 1 (Time Series Observation)
A time series observation, $s^{t}_{i,v}$, refers to a single measurement recorded by sensor $v$ at timestamp $t$ for sample $i$. Sequentially recorded observations over time form a time series $s_{i,v}$ for that sample and sensor.
Definition 2 (Irregular Multivariate Time Series Dataset)
Let $D=\{(S_{i},t_{i},p_{i},y_{i})\}_{i=1}^{N}$ be a dataset where each $S_{i}$ is an irregularly sampled multivariate time series for the $i$-th sample, $t_{i}=\{t_{i,v}\}_{v=1}^{n}$ are the corresponding timestamps, $y_{i}\in\{1,\ldots,C\}$ is the associated categorical outcome label, with $C$ representing the total number of classes, and $p_{i}$ is an optional set of static features (e.g., patient demographics). Each $S_{i}$ consists of multiple time series, one for each sensor $v$. Specifically, $S_{i}=\{s_{i,v}\}_{v=1}^{n}$, where $s_{i,v}$ is the time series of observations recorded by sensor $v$ for sample $i$. The timestamps in $t_{i,v}$ are irregularly spaced, meaning the intervals between consecutive timestamps can vary.
Problem Definition (Irregular Multivariate Time Series Prediction Task) Given a dataset $D=\{(S_{i},t_{i},p_{i},y_{i})\}_{i=1}^{N}$ of irregularly sampled multivariate time series, the goal is to learn a function $f:S_{i}\rightarrow g_{i}$ that maps the sample $s_{i}$, to a fixed-length representation $g_{i}$. This representation can be utilized to predict a label $\hat{y_{i}}\in\{1,\ldots,C\}$ relevant to the downstream task.
4 Methodology
Figure 1 presents the end-to-end pipeline of WaveGNN. In the first step, WaveGNN allocates nodes for each sensor, initializing their states using static features from the dataset, like patient demographics, if available, or randomly if not. Next, each sensor’s irregular observation sequence is embedded using a customized Transformer encoder. This encoder employs masked attention, modified temporal encoding, and a decay mechanism to handle irregularities and capture correlations within each sensor’s sequence. The node states are updated based on these sequence embeddings to represent the latent state of each sensor after the observed window. In the third step, WaveGNN establishes edges that represent the correlations across sensors to effectively utilize inter-series correlations. This allows WaveGNN to compensate for missing observations during the input window by leveraging information from other related sensors to each sensor. These edges are formed based on the updated node states, capturing short-term relationships during the observed window, and on learned global node embeddings, capturing long-term relationships derived from the entire training dataset. Next, WaveGNN performs message passing, which updates the node states to incorporate information from these inferred relationships. This process allows each sensor to benefit from the data captured by other sensors, reducing the impact of irregularity. Finally, the combined node states are aggregated into a single graph-level embedding vector, which is then fed into the prediction head to produce the final prediction. The detailed steps are presented in the following.
4.1 Sensor Graph Construction
Initialization. In WaveGNN, the graph initialization step establishes the preliminary graph structure without connecting any nodes. These will later be updated to reflect the relationships and state of the given sample during the observed time window. The graph comprises nodes, each for a different sensor involved in measuring a time-dependent feature.
For each sample $i$, each node is initialized with a state vector derived from static features $p_{i}$ corresponding to the sample (e.g., demographics such as age and gender and clinical contexts such as ICU types in the medical downstream tasks). If this information is not available for a task, the vectors are initialized randomly. These features are transformed using a two-layer MLP with ReLU activations to create high-dimensional representations that encapsulate the sample’s baseline characteristics and provide a sample-specific initial context for the task:
$Z_{i}=\text{MLP}(p_{i})$ (1)
Here, $p_{i}$ represents the sample $i$’s static features and $Z_{i}=\{z_{i,1},\cdots z_{i,n}\}\in{\mathbb{R}}^{n\times M}$ represents the matrix of initial state vectors for all nodes where $M$ is the embedding dimension.
Observation Embedding. For sample $i$, for each sensor $v$, let the observation at time $t$ be denoted as $s_{i,v}^{t}$. To enhance the expressive power of our model [31], each observation is mapped to a higher-dimensional space using an MLP with $\tanh$ activations:
$h_{i,v}^{t}=\text{MLP}(s_{i,v}^{t})$ (2)
We use $\tanh$ activation to ensure that embeddings are centered around zero, which helps in capturing the inconsistencies in the observations. For instance, two observations that are inconsistent with each other can result in embeddings that are oppositely signed.
Temporal Encoding. To effectively capture the temporal dynamics and patterns within each sensor’s time series, we augment the observation embeddings with time information. Inspired by GRU-D [29], we similarly embed the relative timestamps that reflect the time difference between consecutive observations for each sensor.
To represent these time differences, we apply Time2Vec [32], which encodes the time information into a vector. Time2Vec includes two components: a sinusoidal component that captures periodic patterns, and a non-periodic transformation component that captures non-periodic patterns and time gaps. This dual encoding enables the model to incorporate both regular temporal cycles and irregular time gaps between observations. This turns the time intervals themselves into a valuable source of information rather than merely considering the relative position of observations, as done in the original Transformer [33].
For each observation at time $t_{v,j}$, where $v$ indicates the sensor and $j$ the index of the observation, we compute the time delta $\delta t_{v,j}=t_{v,j}-t_{v,j-1}$ as the difference between the current observation time and the previous observation time for that sensor. The time encoding vector $e_{\delta t_{v,j}}$ is then computed as:
$e_{\delta t_{v,j}}=\text{Time2Vec}(\delta t_{v,j})$ (3)
Augmenting observation embeddings with $e_{\delta t_{v,j}}$ equips the model to better recognize and utilize the underlying temporal structure in the data.
Decay-Based Univariate Sequence Embedding. In order to capture the intra-series correlations, we embed each irregular sequence of observations from each sensor separately. Given the sequence of $T$ observation embeddings $(h_{i,v}^{t_{1}},\cdots,h_{i,v}^{t_{T}})\in\mathbb{R}^{T\times M}$ for sample $i$ and sensor $v$, we add the temporal encoding information to the observation embeddings. For simplicity, we omit the sample index $i$ in the following equations:
$h_{v}^{t_{j}}\leftarrow h_{v}^{t_{j}}+e_{\delta t_{v,j}}\quad\forall j\in\{1,2,\dots,T\}$ (4)
Next, we pass the time-augmented sequence $H_{v}=(h_{v}^{t_{1}},\cdots,h_{v}^{t_{T}})\in\mathbb{R}^{T\times M}$ through a Transformer encoder, using the mask vector $M_{v}=(m_{v,1},\cdots,m_{v,T})\in\mathbb{R}^{T}$ as a padding mask. Here, each $m_{v,j}$ indicates whether there was an observation at time $t_{j}$ for sensor $v$ (1 if present, 0 if missing). If the observation is missing, the corresponding $h_{v}^{t_{j}}$ is a zero-filled place-holder vector. This way, the Transformer’s attention mechanism focuses only on the actual observations, effectively ignoring the missing ones while embedding the real observations:
$Z_{v}=\text{TransformerEncoder}(H_{v},M_{v})$ (5)
The output sequence $Z_{v}=(z_{v}^{t_{1}},\cdots,z_{v}^{t_{T}})$ captures the intra-series correlations, with the attention mechanism ensuring that the presence of missing observations does not negatively affect the representation of the observed data.
Following this, WaveGNN utilizes a decay-based weighting scheme to aggregate the new sequence of embeddings into a single vector. This ensures that more recent observations have a greater influence on the final representation, as they provide the most up-to-date information about the sample. Moreover, WaveGNN’s weighting accounts for the irregular time intervals between observations, considering that observations with shorter intervals between them should have similar impacts on the final embedding, whereas those with larger time gaps should be weighted differently.
To this end, WaveGNN first calculates the time differences of each observation from the most recent observation $t_{v,T}$:
$\delta t^{\prime}_{v,j}=t_{v,T}-t_{v,j}$ (6)
Next, we compute an observation weight at $t_{v,j}$ as follows:
$w_{v,j}=m_{v,j}\cdot e^{-\eta\times\delta t^{\prime}_{v,j}}$ (7)
where $\eta$ is a learned parameter that controls the decay rate. Mask value $m_{v,j}$ is multiplied so the missing observations are ignored. The exponential function is used because it provides a smooth decay of weights controlled by $\eta$ as the time differences increase, ensuring that observations further away from the most recent time have smaller weights.
Finally, we normalize the weights using a softmax function so that they sum up to 1:
$\tilde{w}_{v,j}=\frac{m_{v,j}\cdot e^{-\eta\times\delta t^{\prime}_{v,j}}}{\sum_{k=1}^{T}m_{v,k}\cdot e^{-\eta\times\delta t^{\prime}_{v,k}}}$ (8)
We then aggregate the embeddings to obtain a single vector representation for the sequence using the calculated weights:
$\Delta z_{v}=\sum_{j=1}^{T}\tilde{w}_{v,j}\cdot z_{v}^{t_{j}}$ (9)
The final embedding vector $\Delta z_{v}$ effectively captures the intra-series dependencies and the importance of each observation in sensor $v$. This representation is then used to update the node state vector for the sensor:
$z_{i,v}\leftarrow z_{i,v}+\Delta z_{v}$ (10)
This update is applied in a residual manner to enhance the gradient flow during training and allow the model to refine the initial context derived from static features with information from the actual observations.
Inferring Inter-series Relationships. Now that node states are updated, WaveGNN infers the inter-node relationships by combining two distinct types of adjacency matrices, each representing the relationships from a different perspective:
$\bullet$ Dynamic Short-Term Dependencies ($A_{S}$): This matrix is derived from the multi-head attention mechanism [33] applied to the updated node states. It captures the temporal similarity across node states of different sensors within the observed window, reflecting immediate, short-term dependencies among the nodes. This approach has been shown to be effective in capturing dynamic node relationships [34, 6, 4].
$A_{S}=\text{Multihead\_Attention}(\text{keys}=Z,\text{queries}=Z)$ (11)
$\bullet$ Static Long-Term Similarities ($A_{L}$): WaveGNN learns global node embeddings $\tilde{Z}=\{\tilde{z}_{1},\dots,\tilde{z}_{n}\}\in\mathbb{R}^{n\times M}$, where $M$ is the embedding dimension. These embeddings reflect long-term similarities between sensors, learned from the entire dataset for a specific downstream task. The $A_{L}$ matrix is then computed using the dot product between these global node embeddings, followed by a softmax normalization:
$A_{L}=\text{softmax}(\tilde{Z}\times\tilde{Z}^{T})$ (12)
The final adjacency matrix $A$ is formed by combining the two similarity matrices using a learnable parameter $\alpha$:
$A=\alpha A_{S}+(1-\alpha)A_{L}$ (13)
This allows the model to learn which of these similarities are more important in the downstream task. Finally, we zero out the smallest $K\%$ of the elements in $A$ to make the resulting adjacency matrix sparser.
4.2 Graph-based Prediction
With the updated node states $Z$ and adjacency matrix $A$, the graph $G=(Z,A)$ for the given window of observations is constructed, encapsulating the latest node states and their inter-relationships. We perform GNN-based message passing on this graph to capture the node relationships within the node embeddings. Afterward, we apply pooling and transformations to obtain a single embedding vector representing all important node embeddings. The final embedding is used to make the prediction.
Message Passing. To perform message passing, the graph is passed through a GNN block that utilizes a modified Graph Convolutional Networks (GCN) variant [35]. In this variant, we remove the normalization term and add residual connections between the message-passing layers to preserve directed relationships in the graph and mitigate oversmoothing [36], respectively. This yields the node embeddings matrix $V\in\mathbb{R}^{n\times M^{\prime}}$, where $M^{\prime}$ represents the embedding dimension.
Graph Embedding. To derive a graph-level representation from the node embeddings, two operations are performed: $\bullet$ Max Pooling: Applied to the node embeddings to extract the most significant features, helping to capture the most expressive node attributes relevant to the prediction task.
$\bullet$ Transformation: All node embeddings are concatenated into a single vector, which is then transformed via an MLP to integrate information across all nodes.
Final Prediction. The final $l$-dimensional graph-level representation $g\in\mathbb{R}^{l}$ is obtained by first concatenating the outputs of the max pooling layer and the MLP-transformed concatenated embeddings, followed by an MLP layer for dimensionality reduction.
Finally, the graph-level embedding $g$ is utilized for the downstream task of interest. In our approach, we pass this through another MLP to predict the label of the samples:
$\hat{y}=MLP(g)$ (14)
Where $\hat{y}$ represents the predicted class probabilities.
5 Experiments
5.1 Experimental Setup
Datasets and Evaluation Setup. We evaluate WaveGNN on four real-world healthcare datasets. P12 [37] and MIMIC III [38] contain ICU patient data, with labels indicating survival during hospitalization (48-IHM). MIMIC III also includes labels for multilabel phenotype classification (24-PHE). P19 [39] focuses on sepsis detection, while PAM [40] is an activity monitoring dataset with labels for different activities. P12, P19, and MIMIC III are naturally irregular datasets, while irregularity in PAM is introduced artificially. The dataset statistics’ are presented in Table 1. For a fair comparison, we followed the same preprocessing steps with previous studies [19, 28], and adopted the same dataset splits. To that extent, the training, validation, and testing ratios for P12, P19, and PAM is 80:10:10, and 85:7.5:7.5 for MIMIC III. The indices of these splits are fixed across all methods and experiments. Evaluation metrics also match those used in prior studies to maintain consistency, and all hyperparameters were kept the same across all datasets and experiments. More details regarding the evaluation setup and hyperparameter tuning can be found in the Supplementary Material sections A.1-A.3.
Table 1: Statistics of the datasets used in experimental study.
Statistics MIMIC III P12 P19 PAM
# of samples 16,143 11,988 38,803 5,333
# of timestamps $\leq$ 700 215 60 600
# of sensors 17 36 34 17
static info False True True False
missing ratio (%) 72.07 88.40 94.90 60.00
imbalanced True True True False
# of classes 2 (IHM) / 25 (PHE) 2 2 8
Baselines. We compare WaveGNN to 8 methods: IP-Net [41], DGM2-O [30], mTAND [42] and UTDE [26] are imputation-based baselines, while GRU-D [29], SeFT [20], MTGNN [43], and Raindrop [19], are specialized models designed to handle irregularly sampled datasets. UTDE and Raindrop are considered the state-of-the-art imputation-based and specialized model approaches, respectively.
Table 2: Performance comparison with the state-of-the-art on P12, P19, and PAM datasets. P12 and P19 are evaluated on mortality and sepsis prediction, respectively, based on the AUROC and AUPRC metrics. The PAM dataset is evaluated on activity classification in terms of (weighted) Accuracy, Precision, Recall, and F1 Score. The reported values are averaged across 3 independent runs. Best and second best values are highlighted.
Methods P12 P19 PAM
AUROC AUPRC AUROC AUPRC Accuracy Precision Recall F1 score
GRU-D 81.9 $\pm$ 3.6 46.1 $\pm$ 4.7 83.9 $\pm$ 1.7 46.9 $\pm$ 2.1 83.3 $\pm$ 1.6 84.6 $\pm$ 1.2 85.2 $\pm$ 1.6 84.8 $\pm$ 1.2
SeFT 73.9 $\pm$ 2.5 31.1 $\pm$ 4.1 78.7 $\pm$ 2.4 31.1 $\pm$ 2.8 67.1 $\pm$ 2.2 70.0 $\pm$ 2.4 68.2 $\pm$ 1.5 68.5 $\pm$ 1.8
mTAND 84.2 $\pm$ 0.8 48.2 $\pm$ 3.4 80.4 $\pm$ 1.3 32.4 $\pm$ 1.8 74.6 $\pm$ 4.3 74.3 $\pm$ 4.0 79.5 $\pm$ 2.8 76.8 $\pm$ 3.4
IP-Net 82.6 $\pm$ 1.4 47.6 $\pm$ 3.1 84.6 $\pm$ 1.3 38.1 $\pm$ 3.7 74.3 $\pm$ 3.8 75.6 $\pm$ 2.1 77.9 $\pm$ 2.2 76.6 $\pm$ 2.8
DGM2-O 84.4 $\pm$ 1.6 47.3 $\pm$ 3.1 86.7 $\pm$ 3.4 44.7 $\pm$ 11.7 82.4 $\pm$ 2.3 85.2 $\pm$ 1.2 83.9 $\pm$ 2.3 84.3 $\pm$ 1.8
MTGNN 74.4 $\pm$ 6.7 35.5 $\pm$ 6.0 81.9 $\pm$ 6.2 39.9 $\pm$ 8.9 83.4 $\pm$ 1.9 85.2 $\pm$ 1.7 86.1 $\pm$ 1.9 85.9 $\pm$ 2.4
Raindrop 82.8 $\pm$ 1.7 44.0 $\pm$ 3.0 87.0 $\pm$ 2.3 51.8 $\pm$ 5.5 88.5 $\pm$ 1.5 89.9 $\pm$ 1.5 89.9 $\pm$ 0.6 89.8 $\pm$ 1.0
UTDE 81.1 $\pm$ 2.2 45.9 $\pm$ 2.1 84.3 $\pm$ 3.1 41.1 $\pm$ 6.0 90.4 $\pm$ 1.8 90.7 $\pm$ 1.8 90.4 $\pm$ 1.8 90.4 $\pm$ 1.8
WaveGNN 83.9 $\pm$ 1.2 49.4 $\pm$ 1.5 88.0 $\pm$ 0.9 57.1 $\pm$ 4.7 93.5 $\pm$ 1.1 93.7 $\pm$ 1.1 93.5 $\pm$ 1.1 93.5 $\pm$ 1.1
Improvement -0.59% +2.48% +1.14% +10.23% +3.42% +3.30% +3.42% +3.42%
Table 3: Performance comparison with the state-of-the-art on MIMIC III dataset, for 48-IHM and 24-PHE tasks. The performance of 48-IHM is measured on F1 and AUPRC, and 24-PHE on F1 (Macro) and AUROC. We report the average results across 3 independent runs. Best and second best values are highlighted.
Methods 48-IHM 24-PHE
F1 AUPRC F1 AUROC
GRU-D 42.8 $\pm$ 0.6 45.9 $\pm$ 0.4 19.0 $\pm$ 1.0 73.3 $\pm$ 0.1
SeFT 16.5 $\pm$ 8.6 23.9 $\pm$ 0.5 6.1 $\pm$ 0.2 65.7 $\pm$ 0.1
mTAND 43.9 $\pm$ 0.5 47.5 $\pm$ 1.3 19.9 $\pm$ 0.4 73.5 $\pm$ 0.1
IP-Net 37.2 $\pm$ 2.8 39.4 $\pm$ 1.1 17.9 $\pm$ 0.7 73.5 $\pm$ 0.1
DGM2-O 39.1 $\pm$ 1.5 37.8 $\pm$ 1.5 18.4 $\pm$ 0.2 71.7 $\pm$ 0.2
MTGNN 38.6 $\pm$ 2.5 36.5 $\pm$ 2.1 14.5 $\pm$ 1.7 70.6 $\pm$ 0.7
Raindrop 39.5 $\pm$ 3.7 36.2 $\pm$ 0.4 21.8 $\pm$ 1.7 74.0 $\pm$ 0.9
UTDE 45.3 $\pm$ 0.7 49.6 $\pm$ 1.0 24.9 $\pm$ 0.4 75.6 $\pm$ 0.2
WaveGNN 46.5 $\pm$ 0.9 47.8 $\pm$ 1.3 30.8 $\pm$ 0.2 74.9 $\pm$ 0.1
Improvement +2.64% -3.62% +23.69% -0.93%
5.2 Results
Overall Performance. We report the overall performance of WaveGNN and baselines in Tables 2 and 3, where the best and second best values are highlighted. WaveGNN consistently outperforms the baselines in the majority of cases (9 out of 12 experiments), and when it doesn’t, its performance is very close to the best-performing baseline.
Notably, WaveGNN achieves significant improvements in AUPRC, with gains of 10.2% and 2.4% over the best baseline in the binary tasks on the P19 and P12 datasets, respectively. In the multi-class task on the PAM dataset, WaveGNN consistently outperforms all baselines, demonstrating performance improvements of over 3.3% across all metrics. The PAM dataset has the smallest missing ratio and very long sequences among the evaluated datasets. This highlights that WaveGNN is highly effective not only in challenging scenarios like P19, which has over a 94% missing ratio and short sequences, but also in cases with less irregularity.
The results on the binary 48-IHM task of the MIMIC dataset are mixed. While WaveGNN outperforms UTDE by 2.64% in terms of F1 score, UTDE performs better in terms of AUPRC. This can be attributed to MIMIC’s longer sequences of observations with a lower missing data ratio compared to the P12 and P19 binary tasks, making imputation-based methods more effective for this less challenging binary classification task. However, in the more challenging multi-label classification task of 24-PHE on the same dataset, WaveGNN outperforms UTDE by a margin of 23.7% in F1 score, demonstrating that imputation-based methods become less effective in more complex scenarios.
In sum, on datasets with lower missing data ratios and longer sequences, like MIMIC and PAM, imputation-based methods like UTDE tend to outperform specialized models. Conversely, on datasets with higher missing data ratios, such as P19, specialized models like Raindrop are more effective. This makes sense: with more observations, imputation can more accurately estimate missing values, whereas in cases with fewer observations, the biased assumptions of imputation methods become less effective, making direct modeling approaches more suitable. Notably, WaveGNN consistently outperforms the best of both methods across all datasets.
Leave-sensors-out settings. We also evaluate WaveGNN under more challenging scenarios where whole sensor observations are missing. This simulates real-world cases where some sensors fail or are unavailable. Following the approach adopted by Raindrop [19], we experimented with two setups on PAM dataset. In the first one, we remove all the observations for a fixed set of the most informative sensors from samples, while in the second setting, we remove all the observations for a random set of sensors. Table 4 presents the results for both cases when 10%, 30%, and 50% of the sensors are removed. Our approach outperforms the baselines in 23 out of 24 experiments and ranks second in only one case. Specifically, WaveGNN demonstrates notable improvements, with up to 15.5% relative improvement in F1 score compared to the best baseline with 10% missing sensors, and up to 45.2% and 7.5% improvement for cases with 30% and 50% missing sensors, respectively.
Among the baselines, Raindrop remains the strongest competitor to WaveGNN. This shows the importance of capturing inter-series relationships and highlights the effectiveness of graph-based modeling. By successfully capturing the dependencies between sensors, missing data can be compensated, leveraging insights from correlated sensors that are relevant to the final outcome. In contrast, imputation-based methods, like mTAND and UTDE, demonstrate the largest performance degradation as the missing sensor ratio increases. These approaches still try to fill in the missing values, but with entire sensor observations missing, they have little to no reliable data to base their imputations on. This leads to inaccurate or biased reconstructions and, hence, affects their performance.
Table 4: Performance comparison on the PAM dataset under different missing sensor ratios and different settings. We report the average results across 3 independent runs. Best and second best values are highlighted.
Experiment Setting Leave-fixed-sensors-out Leave-random-sensors-out
Missing sensor ratio Methods Accuracy Precision Recall F1 score Accuracy Precision Recall F1 score
10% GRU-D 65.4 $\pm$ 1.7 72.6 $\pm$ 2.6 64.3 $\pm$ 5.3 63.6 $\pm$ 0.4 68.4 $\pm$ 3.7 74.2 $\pm$ 3.0 70.8 $\pm$ 4.2 72.0 $\pm$ 3.7
SeFT 58.9 $\pm$ 2.3 62.5 $\pm$ 1.8 59.6 $\pm$ 2.6 59.6 $\pm$ 2.6 40.0 $\pm$ 1.9 40.8 $\pm$ 3.2 41.0 $\pm$ 0.7 39.9 $\pm$ 1.5
mTAND 58.8 $\pm$ 2.7 59.5 $\pm$ 5.3 64.4 $\pm$ 2.9 61.8 $\pm$ 4.1 53.4 $\pm$ 2.0 54.8 $\pm$ 2.7 57.0 $\pm$ 1.9 55.9 $\pm$ 2.2
RAINDROP 77.2 $\pm$ 2.1 82.3 $\pm$ 1.1 78.4 $\pm$ 1.9 75.2 $\pm$ 3.1 76.7 $\pm$ 1.8 79.9 $\pm$ 1.7 77.9 $\pm$ 2.3 78.6 $\pm$ 1.8
UTDE 71.8 $\pm$ 1.7 76.5 $\pm$ 2.6 71.8 $\pm$ 1.7 70.6 $\pm$ 2.3 74.5 $\pm$ 1.9 75.9 $\pm$ 2.1 74.5 $\pm$ 1.9 74.6 $\pm$ 2.1
WaveGNN 86.9 $\pm$ 2.2 88.0 $\pm$ 1.3 86.9 $\pm$ 2.2 86.9 $\pm$ 3.3 82.2 $\pm$ 1.5 83.9 $\pm$ 0.8 82.2 $\pm$ 1.5 82.5 $\pm$ 1.3
Improvement +12.56% +6.92% +10.84% +15.55% +7.17% +5.00% +5.51% +4.96%
30% GRU-D 45.1 $\pm$ 2.9 51.7 $\pm$ 6.2 42.1 $\pm$ 6.6 47.2 $\pm$ 3.9 58.0 $\pm$ 2.0 63.2 $\pm$ 1.7 58.2 $\pm$ 3.1 59.3 $\pm$ 3.5
SeFT 32.7 $\pm$ 2.3 27.9 $\pm$ 2.4 34.5 $\pm$ 3.0 28.0 $\pm$ 1.4 31.7 $\pm$ 1.5 31.0 $\pm$ 2.7 32.0 $\pm$ 1.2 28.0 $\pm$ 1.6
mTAND 27.5 $\pm$ 4.5 31.2 $\pm$ 7.3 30.6 $\pm$ 4.0 30.8 $\pm$ 5.6 34.7 $\pm$ 5.5 43.4 $\pm$ 4.0 36.3 $\pm$ 4.7 39.5 $\pm$ 4.4
RAINDROP 52.4 $\pm$ 2.8 60.9 $\pm$ 3.8 51.3 $\pm$ 1.6 48.4 $\pm$ 5.6 60.3 $\pm$ 3.3 68.1 $\pm$ 4.6 60.3 $\pm$ 3.3 60.3 $\pm$ 3.3
UTDE 41.2 $\pm$ 2.9 46.9 $\pm$ 6.5 41.2 $\pm$ 2.9 35.7 $\pm$ 3.9 57.9 $\pm$ 2.5 65.0 $\pm$ 3.4 57.9 $\pm$ 2.5 58.4 $\pm$ 2.8
WaveGNN 71.2 $\pm$ 2.8 75.7 $\pm$ 3.4 71.2 $\pm$ 2.8 70.3 $\pm$ 3.2 68.6 $\pm$ 1.6 75.1 $\pm$ 1.8 68.6 $\pm$ 1.6 69.0 $\pm$ 1.3
Improvement +35.87% +24.30% +38.79% +45.24% +13.76% +10.27% +13.76% +14.42%
50% GRU-D 37.3 $\pm$ 2.7 29.6 $\pm$ 5.9 32.8 $\pm$ 4.6 26.6 $\pm$ 5.9 49.7 $\pm$ 1.2 52.4 $\pm$ 0.3 42.5 $\pm$ 1.7 47.5 $\pm$ 1.2
SeFT 24.7 $\pm$ 1.1 15.9 $\pm$ 2.7 25.3 $\pm$ 2.6 18.2 $\pm$ 2.4 26.4 $\pm$ 1.4 23.0 $\pm$ 2.9 27.5 $\pm$ 0.4 23.5 $\pm$ 1.8
mTAND 16.9 $\pm$ 3.1 12.6 $\pm$ 5.5 17.0 $\pm$ 1.6 13.9 $\pm$ 4.0 20.9 $\pm$ 3.1 35.1 $\pm$ 6.1 23.0 $\pm$ 3.2 27.7 $\pm$ 3.9
RAINDROP 46.6 $\pm$ 2.6 44.5 $\pm$ 2.6 42.4 $\pm$ 3.9 38.0 $\pm$ 4.0 47.2 $\pm$ 4.4 59.4 $\pm$ 3.9 44.8 $\pm$ 3.5 47.6 $\pm$ 5.2
UTDE 30.6 $\pm$ 2.4 32.9 $\pm$ 13.5 30.6 $\pm$ 2.4 25.1 $\pm$ 3.9 43.5 $\pm$ 1.7 52.7 $\pm$ 4.0 43.5 $\pm$ 1.7 43.2 $\pm$ 1.9
WaveGNN 43.2 $\pm$ 3.5 46.6 $\pm$ 6.1 43.2 $\pm$ 3.5 38.2 $\pm$ 3.4 51.2$\pm$ 3.3 63.3 $\pm$ 5.1 51.2 $\pm$ 3.3 51.2 $\pm$ 3.3
Improvement -7.29% +4.71% +1.88% +0.52% +3.01% +6.56% +14.28% +7.56%
5.3 Ablation Study
We created different variations of WaveGNN to understand the effectiveness of its different components: w/o short-term and w/o long-term variants remove the $A_{S}$ and $A_{L}$ similarity matrices, so that the short- and long-term inter-series relationships are not considered, respectively. The w/o inter-series variant excludes the graph message passing part, so neither of the inter-series similarity measures are considered, while the w/o intra-series variant removes the transformer-based module and simply aggregates the observation embeddings for each sensor to obtain a univariate sequence embedding. Finally, the w/o temporal encoding variant discards the temporal encoding module and uses a simple positional encoding for the Transformer, and the w/o decay rate variant removes WaveGNN’s decay mechanism, aggregating the output sequence embedding of the Transformer with equal weights, thus treating all observations in a sequence equally.
Table 5 presents the results of the ablation study on the PAM dataset, showing the impact of removing various components of WaveGNN. Our findings indicate that all components contribute to improved performance, with the most significant drops occurring when inter-series and intra-series relationships are removed, leading to performance decreases of 33.8% and 13.2% in weighted F1 score, respectively. This shows the importance of simultaneously addressing both intra-variable dynamics and inter-variable dependencies to effectively model irregular datasets.
Interestingly, the removal of dynamic short-term dependencies has a greater negative impact on performance than the removal of long-term dependencies. This outcome is intuitive, as short-term dependencies between sensors within the current input window are often more critical than long-term dependencies, given their specificity to the individual sample and time window, rather than being generalized across all samples and time periods. For example, the relationship between a patient’s heart rate and blood sugar level during a particular time window might differ from the typical pattern usually seen in that patient, possibly due to the effects of medication or other factors, making it more relevant to focus on these short-term dependencies.
Lastly, both the temporal encoding and the decay mechanism components are integral to WaveGNN’s overall performance. Temporal encoding allows the model to capture the temporal dynamics and patterns within each sensor’s time series, while the decay mechanism ensures that more recent data, which is often more indicative of final outcomes, is given appropriate weight. Together, these components enable WaveGNN to manage the timing and relevance of observations effectively, contributing to overall performance.
Table 5: Ablation study results on the PAM dataset. We report the average results across 3 independent runs.
Ablation Setting Accuracy Precision Recall F1 score
w/o short-term 89.9 $\pm$ 1.9 90.1 $\pm$ 1.9 89.9 $\pm$ 1.9 89.8 $\pm$ 2.0
w/o long-term 92.3 $\pm$ 1.9 92.4 $\pm$ 1.9 92.3 $\pm$ 1.9 92.2 $\pm$ 1.9
w/o inter-series 64.1 $\pm$ 1.7 63.5 $\pm$ 2.0 64.1 $\pm$ 3.1 61.9 $\pm$ 3.1
w/o intra-series 81.2 $\pm$ 1.3 81.5 $\pm$ 1.5 81.2 $\pm$ 1.3 81.2 $\pm$ 1.4
w/o temporal encoding 90.6 $\pm$ 1.1 90.6 $\pm$ 1.0 90.6 $\pm$ 1.1 90.4 $\pm$ 1.2
w/o decay rate 90.6 $\pm$ 2.3 90.7 $\pm$ 2.3 90.6 $\pm$ 2.3 90.5 $\pm$ 2.3
WaveGNN 93.5 $\pm$ 1.1 93.7 $\pm$ 1.1 93.5 $\pm$ 1.1 93.5 $\pm$ 1.1
6 Conclusion
In this study, we introduced WaveGNN, a novel framework designed to directly embed irregularly sampled multivariate time series datasets for accurate predictions. Unlike previous imputation-based and specialized model approaches, WaveGNN simultaneously captures both intra-series correlations and inter-series relationships, enabling the model to effectively compensate for missing observations. Our evaluation demonstrated that WaveGNN outperforms state-of-the-art baselines across diverse datasets, even under conditions of extreme sparsity where entire sensors are missing, thereby validating the importance of each proposed component within the framework. In the future, we plan to expand WaveGNN to support multimodal scenarios.